OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. P(6, 2), Q(8, -4)

OpenStudy (anonymous):

When two points are given as \((x_1, y_1)\) and \((x_2,y_2)\), then two point slope form is given as: \[\frac{y - y_1}{y_2-y_1} = \frac{x - x_1}{x_2 - x_1}\]

OpenStudy (anonymous):

Just plug in the values and then we will solve it further..

OpenStudy (anonymous):

okay so this how ii plug in the the number y-2 x-6 ---- = ----- -4-2 -4-6